# TVR-Ranking: A Dataset for Ranked Video Moment Retrieval with Imprecise Queries

### v1

@@ 1 Introduction
Given a query expressed in natural language, to retrieve or to locate a temporal moment from video(s) that semantically matches the query has many applications. A temporal moment refers to a segment within a source video with identified start and end timestamps. Examples of such applications include searching for a specific scene in security surveillance videos Yuan et al. (2023), locating a medical procedure within an educational tutorial Gupta et al. (2023), or identifying desired scenes for video editing purposes, among others.
A few tasks with different names have been studied for addressing similar objectives, including video retrieval (VR), video moment retrieval (VMR), natural language video localization (NLVL), temporal sentence grounding in video (TSGV), and video corpus moment retrieval (VCMR) Liu et al. (2023); Zhang et al. (2023). Among them, VR involves retrieving a video from a collection based on visual content, akin to video search on platforms like YouTube, but with the search criteria grounded in the visual content of videos. NLVL and TSGV, more commonly used in CV and NLP communities, refer to the same task as video moment retrieval (VMR) in the IR community. VMR aims to locate a moment within a given video that semantically matches the text query. The VCMR task is a direct extension of VMR, focusing on retrieving a moment from a collection of videos Victor et al. (2019). As depicted in Figure 1, existing tasks VR VMR and VCMR, all aim to find one answer, being either a video or a moment, for a given query.
@@
The reason for expecting one exact answer to a query lies primarily in the annotation of benchmark datasets. During annotation, annotators watch a video, then provide textual descriptions of meaningful video moments in this video. Subsequently, each description serves as the query to retrieve the corresponding moment from this source video. Given that a query typically describes a specific moment precisely, a model trained on these datasets can assume the existence of the moment to be searched for, and all queries are from users who possess a good understanding of the source video.
@@
To date, no datasets cater to this novel task setting. Thus, we have curated the TVR-Ranking dataset to facilitate the RVMR task. As its name suggests, the dataset has its root in the TVR  [^1] The TVR dataset is under the MIT License. dataset Lei et al. (2020). Specifically, we reuse the source videos in TVR, i.e., video clips from six TV series, and the original moment annotations i.e., the begin/end timestamps of meaningful moments. We have made two main efforts in the development of the TVR-Ranking. The first is deriving imprecise queries from the original moment descriptions. In the TVR dataset, moment descriptions are very detailed and often contain TV character names. We substitute character names with pronouns using carefully crafted prompts to ChatGPT, with follow up quality control. This process entails replacing a total of 160,701 words across 72,842 moment descriptions. Our second effort involves annotating the relevance scores of moments to these imprecise queries. This task was undertaken by 23 annotators over 1,200 working hours. We have annotated relevant moments for 3,281 queries. Among them, 52% of queries were each annotated with 20 candidate moments, featuring five relevance levels ranging from irrelevant (0) to a perfect match (4). The remaining 48% of queries were each annotated with 40 candidate moments. The annotation of the relevance score for a candidate moment to a query (i.e., a query-moment pair) was conducted by either two annotators (in case of consensus can be reached by the two) or four annotators. A total of 94,442 relevance scores were annotated with consensus. The annotated queries are divided into 500 validation and 2,781 test queries.
@@ 2 Related work
Charads-STA Gao et al. (2017) 16,128 11,770 6,672 1,303 469 26.33 56.69
ActivityNet Captions Krishna et al. (2017) 54,559 54,559 14,926 13,645 4,510 560.37 487.60
TACoS  Regneri et al. (2013) 18,227 7,069 127 2,287 994 2220.82 10.11
MAD (v2-unnamed)  Soldan et al. (2022) 3,328,745 328,742 488 56,066 13,993 263.46 1,207.3
Ego4D(NLQ) Grauman et al. (2022) 18,399 18,374 1,685 3,337 823 56.85 231.82
UCA Yuan et al. (2023) 19,211 18,299 1,544 4,087 1,580 84.39 96.71
MedVidQA Gupta et al. (2022) 3,010 2,990 899 2,291 670 51.78 95.72
QVHighlight Lei et al. (2021) 10,310 18,367 10,148 7,750 1,824 125.50 422.83
TVR Lei et al. (2020) 98,070 97,442 19,614 18,856 6,104 240.70 414.86
@@
The first assumption stems from traditional annotation processes where annotators are required to watch the entire video and then describe meaningful moments therein. Most datasets listed in Table 1 are annotated in this way, including DiDeMo Anne Hendricks et al. (2017), TACoS Regneri et al. (2013), TVR Lei et al. (2020), ActivityNet Caption Krishna et al. (2017), Ego4D(NLQ) Grauman et al. (2022), and UCA Yuan et al. (2023). Besides, the Charades-STA dataset Gao et al. (2017) extends the Charades dataset Sigurdsson et al. (2016) by segmenting video descriptions into sentences and linking them to specific video timestamps via keywords. In contrast, queries in our dataset may or may not provide precise descriptions of moments, thus embracing users with different levels of understanding of the corpus.
Typically, standard VMR datasets generally link a query to a single relevant moment. For instance, the health-related queries in the MedVidQA dataset Gupta et al. (2022), sourced from WikiHow’s ‘Health’ category, are well-represented of real-world scenarios. Nonetheless, this dataset confines each query to just the most relevant moment. In contrast, real-life situations frequently encompass multiple moments that can be similarly described. The QVHighlight dataset Lei et al. (2021) was pioneering in allowing queries to match multiple moments within a single video. However, it still restricts searches to single videos and focuses only on perfect matches. Our dataset aims to retrieve a ranked list of moments from a video corpus based on imprecise queries, functioning more like a search engine and accommodating both closely and loosely relevant matches. This paradigm not only broadens the utility of the results but also aligns more closely with practical search needs.
@@ 3 The TVR-Ranking Dataset Annotation
In the ideal setting, a dataset shall well reflect the context of real-world applications, e.g., the data source and the information needs from users Yu and Sun (2023). In the RVMR task setting, we assume there exists a collection of videos, and users search for relevant moments through textual descriptions as queries. However, such kinds of queries can only be collected from logs of video search services, which are not publicly accessible. Without access to such resources, we choose to derive user queries from existing data annotations, i.e., the datasets listed in Table 1.
Our immediate task is to choose which existing dataset to use as the raw data for annotation. To this end, we compare the existing datasets based on the following perspectives: accessibility to the raw videos, number of videos, variants of different activities/scenes, and number of moments annotated. Because existing annotations are mostly descriptions of scenes and/or actions, the number of verbs has been widely used to measure the number of activities covered in a dataset Sigurdsson et al. (2017). Based on these considerations, we adopt the TVR dataset as the raw source for our annotation; accordingly, our dataset is named TVR-Ranking.
The TVR dataset contains video clips from six different TV series, along with their corresponding subtitles and audio tracks. In the construction of the original TVR dataset, annotators were tasked with identifying the boundaries of events i.e., moments, within these clips, and describing their content. Workers also provided whether the description was purely based on the visual content, the subtitle, or both the video and subtitle. The TVR dataset comprises 72,842 video-only, 8,920 subtitle-only, and 16,308 video-subtitle moment descriptions. [^2] The numbers reported here are from the dataset version used in our annotation, with negligible differences from those reported in the original paper Lei et al. (2020). In our TRV-Ranking construction, we aim to concentrate on the visual aspects; therefore, we only consider the annotations that are purely based on the visual content of the videos.
@@ 3.1 Imprecise Queries
The character name substitution is through carefully designed prompts to ChatGPT (see Appendix A), with quality checks. The output of ChatGPT is a quality substitution if it successfully passes two validation checks. The first check is for semantic consistency, to ensure no significant change in terms of semantic meaning after substitution. The SimCSE Gao et al. (2021) similarity of moment descriptions before and after the substitution is expected to be above a threshold (0.4 in our implementation). The second check is to ensure no person names appear in the substituted version. We detect person names in the substituted moment description using Flair Akbik et al. (2019). If a substitution fails to pass both checks, the moment description undergoes human review and is fed to ChatGPT again for another substitution with a different temperature parameter setting, till it passes the two checks.
The above procedure replaces 160,701 words across 72,842 moment descriptions, averaging 2.21 words per description. Table 3 presents the top 12 most frequently replaced words, which shows a diversity of personal pronouns in the descriptions.333 The implementation details and the prompt are detailed in Appendix A.
@@ 3.2 Relevance Annotation and Quality Control
Let $m.c$ and $m.v$ represent the caption and the visual content of moment $m$, respectively. To annotate the ground truth moments for a query $q$, we initially retrieve the top-$K$ moment candidates based on their similarity to $q$ by using SimCSE Gao et al. (2021), denoted by using $sim(q,m.c)$. In our annotation process, we set $K=20$ for the first batch. This batch of 20 query-moment pairs is then presented to two annotators. [^4] The query $q$ is the same in the batch of 20 query-moment pairs $\langle q,m.v\rangle$. However, during annotation, an annotator is presented with one query-moment pair each time through the annotation interface. Each annotator independently labels the degree of relevance of every query-moment pair $\langle q,m.v\rangle$ purely based on the moment’s visual content, assigning a score from 0 for irrelevant, to 4 for a perfect match.
If the difference between the two relevance scores assigned by two annotators is either 1 or 0, then we considered the two annotators to have reached a consensus. The average relevance score was then rounded up to the nearest whole number as the final score for this query-moment pair. If the two annotators fail to reach a consensus, the same query-moment pair is assigned to another two annotators. Then we have a total of 4 scores. Among the 4, we remove one highest score and one lowest relevance score. If the difference between the remaining two scores is either 1 or 0, we consider a consensus is reached; the average of the remaining two scores is rounded up to the nearest whole number as the final score. Otherwise, the pair is discarded.
@@ 4 Evaluation Metric for RVMR
NDCG for Ranked Retrieval. Normalized Discounted Cumulative Gain (NDCG) is delicately designed for evaluating ranking results with different relevance levels Manning et al. (2008). Specifically, the Discounted Cumulative Gain (DCG) of the top $K$ ranked results is defined in Equation 1, where $i$ is the ranking position with 1 being the top ranked position, $rel_{i}$ is the level of relevance. For example, the left part of Figure 2(b) shows four ground truth moments with $g_{1}$ at rank 1 position and $g_{4}$ at rank 4 position. To their left are the relevance levels with $rel_{1}=4$ for $g_{1}$ and $rel_{2}=2$ for $g_{2}$.
@@
Figure 2 shows the computation of $NDCG@3,IoU\geq 0.3$ as an example. Starting with top predicted moment $p_{1}$, where $IoU(p_{1},g_{1})=0.35$ and $IoU(p_{1},g_{3})=0.4$, both exceeding the threshold $\mu=0.3$, as shown in Figure 2(b). Since the $IoU$ with $g_{3}$ is higher, we consider $p_{1}$ matching with $g_{3}$ and assign $g_{3}$’s relevance score to $p_{1}$. Then $g_{3}$ is removed from the ground truth due to being matched by a predicted moment. Assuming $p_{2}$ is very similar to $p_{1}$ (or a new duplicate), [^5] This is a rare case, and it is unlikely to have two ground truth moments matching duplicate predictions as well. However, we would like to show that our proposed measure is able to handle such a rare case. with $IoU(p_{2},g_{1})=0.35$ and $IoU(p_{2},g_{3})=0.4$, as depicted in Figure 2(c). Since $g_{3}$ has been removed from the ground truth, $g_{1}$ becomes the sole match for $p_{2}$, resulting in a relevance score of 4 for $p_{2}$, and the removal of $g_{1}$ from the ground truth. For $p_{3}$, $IoU(p_{3},g_{4})=0.5$, indicating a match with $g_{4}$, thus $p_{3}$ is assigned a relevance score of 2, and $g_{4}$ is removed. The relevance scores of the predicted moments are 2, 4, and 2. To compute $NDCG$ of the ground truth ranking, we consider a perfect ranking of the top 3 ground truth moments (4, 2, and 2), regardless of whether they are matched by predicted moments or not, as illustrated by $K=3$ on the left side of Figure 2(b).
@@ 5 Baseline Performance
Illustrated in Figure 1, the closest task setting to RVMR is VCMR. In particular, if a VCMR model can compute a form of confidence for its retrieval result, then a ranking of the predicted moments can be easily achieved. Hence, we adapt three VCMR models to RVMR and evaluate them on the TRV-Ranking :XML Lei et al. (2020), CONQUER Hou et al. (2021), and ReLoCLNet Zhang et al. (2021). The main change in the adaptations is the introduction of weight to the training loss based on query-moment similarity i.e., $sim(q,m.c)$, recognizing that moments vary in relevance to a query. This weight aims to diminish the influence of less relevant moments on model training by adjusting the loss. The implementation details are in Appendix C.1.
@@
XML Lei et al. (2020) 1 0.1015 0.0998 0.0751 0.0731 0.0289 0.0313
20 0.1670 0.1287 0.0530 0.1554 0.1155 0.0478
40 0.1767 0.1356 0.0582 0.1624 0.1228 0.0514
CONQUER Hou et al. (2021) 1 0.1143 0.1029 0.0956 0.0811 0.0633 0.0539
20 0.2014 0.1946 0.1817 0.1777 0.1363 0.1313
40 0.2068 0.1919 0.1875 0.1768 0.1428 0.1291
ReLoCLNet Zhang et al. (2021) 1 0.1234 0.1236 0.1110 0.1100 0.0712 0.0770
20 0.2615 0.2545 0.2424 0.2373 0.1752 0.1762
40 0.2715 0.2700 0.2552 0.2541 0.1850 0.1913
We conducted three sets of experiments with different $N=\{1,20,40\}$ values for a comprehensive evaluation. We adapted the evaluation metric to $NDCG@K,IoU\geq\mu$, for $K=\{10,20,40\}$ and $\mu=\{0.3,0.5,0.7\}$. The parameter search is based on the best results on the validation set with $NDCG@20,IoU\geq 0.5$. Table 5 reports the experiment results with $NDCG@20$ and the full results are in Appendix C.
@@
The $K$ Values for $NDCG$. Across various $K$ values, we observe no significant differences in their respective performances when other settings, such as the model, $K$, and $\mu$, are held constant. If a model is good at result ranking, its performance in ranking the top 10 results should be consistent with its performance in ranking the top 20 results.
@@
The Choice of $N$ in Pseudo Training: Across different values of $N$, the trend is consistent among the models tested. The lowest performances are consistently observed at $N=1$, while there are no significant differences between $N=20$ and $N=40$, suggesting that $N=20$ adequately covers a wide range of relevant moments.
Comparison of Baseline Models: All three models (XML, CONQUER, and ReLoCNet) exhibit consistent performance trends across different training sets and metrics. Among the three, ReLoCNet emerges as the top performer, notably when sufficient training moments are provided i.e., $N=20$ or $N=40$. However, the three baselines show different performance ranking on the VCMR task, where CONQUER demonstrates superior performance against ReLoCNet, and XML is slightly better than ReLoCNet as well. The discrepancy on RVMR implies that the abilities required by our RVMR task differ from those of the VCMR task.

### v2

@@ 1 Introduction
Given a query expressed in natural language, to retrieve or to locate a temporal moment from video(s) that semantically matches the query has many applications. A temporal moment refers to a segment within a source video with identified start and end timestamps. Examples of such applications include searching for a specific scene in security surveillance videos [1], locating a medical procedure within an educational tutorial [2], or identifying desired scenes for video editing purposes, among others.
A few tasks with different names have been studied for addressing similar objectives, including video retrieval (VR), video moment retrieval (VMR), natural language video localization (NLVL), temporal sentence grounding in video (TSGV), and video corpus moment retrieval (VCMR) [3, 4]. Among them, VR involves retrieving a video from a collection based on visual content, akin to video search on platforms like YouTube, but with the search criteria grounded in the visual content of videos. NLVL and TSGV, more commonly used in CV and NLP communities, refer to the same task as video moment retrieval (VMR) in the IR community. VMR aims to locate a moment within a given video that semantically matches the text query. The VCMR task is a direct extension of VMR, focusing on retrieving a moment from a collection of videos [5]. As depicted in Figure 1, existing tasks VR VMR and VCMR, all aim to find one answer, being either a video or a moment, for a given query.
@@
The reason for expecting one exact answer to a query in existing datasets lies primarily in the annotation of benchmark datasets. During annotation, annotators watch a video, then provide textual descriptions of meaningful video moments in this video. Subsequently, each description serves as the query to retrieve the corresponding moment from this source video. Given that a query typically describes a specific moment precisely, a model trained on these datasets can assume the existence of the moment to be searched for, and all queries are from users who possess a good understanding of the source video.
@@
To date, no datasets cater to this novel task setting. Thus, we have curated the TVR-Ranking dataset to facilitate the RVMR task. As its name suggests, the dataset has its root in the TVR  [^1] The TVR dataset is under the MIT License. dataset [6]. Specifically, we reuse the source videos in TVR, i.e., video clips from six TV series, and the original moment annotations i.e., the begin/end timestamps of meaningful moments. We have made two main efforts in the development of the TVR-Ranking. The first is deriving imprecise queries from the original moment descriptions. In the TVR dataset, moment descriptions are very detailed and often contain TV character names. We substitute character names with pronouns using carefully crafted prompts to ChatGPT, with follow up quality control. This process entails replacing a total of 160,701 words across 72,842 moment descriptions. Our second effort involves annotating the relevance scores of moments to these imprecise queries. This task was undertaken by 23 annotators over 1,200 working hours. We have annotated relevant moments for 3,281 queries. Among them, 52% of queries were each annotated with 20 candidate moments, featuring five relevance levels ranging from irrelevant (0) to a perfect match (4). The remaining 48% of queries were each annotated with 40 candidate moments. The annotation of the relevance score for a candidate moment to a query (i.e., a query-moment pair) was conducted by either two annotators (in case of consensus can be reached by the two) or four annotators. A total of 94,442 relevance scores were annotated with consensus. The annotated queries are divided into 500 validation and 2,781 test queries.
@@ 2 Related work
Charads-STA [7] 16,128 11,770 6,672 1,303 469 26.33 56.69
ActivityNet Captions [8] 54,559 54,559 14,926 13,645 4,510 560.37 487.60
TACoS  [9] 18,227 7,069 127 2,287 994 2220.82 10.11
MAD (v2-unnamed) [10] 3,328,745 328,742 488 56,066 13,993 263.46 1,207.3
Ego4D(NLQ) [11] 18,399 18,374 1,685 3,337 823 56.85 231.82
UCA [1] 19,211 18,299 1,544 4,087 1,580 84.39 96.71
MedVidQA [12] 3,010 2,990 899 2,291 670 51.78 95.72
QVHighlight [13] 10,310 18,367 10,148 7,750 1,824 125.50 422.83
TVR [6] 98,070 97,442 19,614 18,856 6,104 240.70 414.86
@@
The first assumption stems from traditional annotation processes where annotators are required to watch the entire video and then describe meaningful moments therein. Most datasets listed in Table 1 are annotated in this way, including DiDeMo [14], TACoS [9], TVR [6], ActivityNet Caption [8], Ego4D(NLQ) [11], and UCA [1]. Besides, the Charades-STA dataset [7] extends the Charades dataset [15] by segmenting video descriptions into sentences and linking them to specific video timestamps via keywords. In contrast, queries in our dataset may or may not provide precise descriptions of moments, thus embracing users with different levels of understanding of the corpus.
Typically, standard VMR datasets generally link a query to a single relevant moment. For instance, the health-related queries in the MedVidQA dataset [12], sourced from WikiHow’s ‘Health’ category, are well-represented of real-world scenarios. Nonetheless, this dataset confines each query to just the most relevant moment. In contrast, real-life situations frequently encompass multiple moments that can be similarly described. The QVHighlight dataset [13] was pioneering in allowing queries to match multiple moments within a single video. However, it still restricts searches to single videos and focuses only on perfect matches. Our dataset aims to retrieve a ranked list of moments from a video corpus based on imprecise queries, functioning more like a search engine and accommodating both closely and loosely relevant matches. This paradigm not only broadens the utility of the results but also aligns more closely with practical search needs.
@@ 3 The TVR-Ranking Dataset Annotation
In the ideal setting, a dataset shall well reflect the context of real-world applications, e.g., the data source and the information needs from users [16]. In the RVMR task setting, we assume there exists a collection of videos, and users search for relevant moments through textual descriptions as queries. However, such kinds of queries can only be collected from logs of video search services, which are not publicly accessible. Without access to such resources, we choose to derive user queries from existing data annotations, i.e., the datasets listed in Table 1.
Our immediate task is to choose which existing dataset to use as the raw data for annotation. To this end, we compare the existing datasets based on the following perspectives: accessibility to the raw videos, number of videos, variants of different activities/scenes, and number of moments annotated. Because existing annotations are mostly descriptions of scenes and/or actions, the number of verbs has been widely used to measure the number of activities covered in a dataset [17]. Based on these considerations, we adopt the TVR dataset as the raw source for our annotation; accordingly, our dataset is named TVR-Ranking.
The TVR dataset contains video clips from six different TV series, along with their corresponding subtitles and audio tracks. In the construction of the original TVR dataset, annotators were tasked with identifying the boundaries of events i.e., moments, within these clips, and describing their content. Workers also provided whether the description was purely based on the visual content, the subtitle, or both the video and subtitle. The TVR dataset comprises 72,842 video-only, 8,920 subtitle-only, and 16,308 video-subtitle moment descriptions. [^2] The numbers reported here are from the dataset version used in our annotation, with negligible differences from those reported in the original paper [6]. In our TRV-Ranking construction, we aim to concentrate on the visual aspects; therefore, we only consider the annotations that are purely based on the visual content of the videos.
@@ 3.1 Imprecise Queries
The character name substitution is through carefully designed prompts to ChatGPT (see Appendix A), with quality checks. The output of ChatGPT is a quality substitution if it successfully passes two validation checks. The first check is for semantic consistency, to ensure no significant change in terms of semantic meaning after substitution. The SimCSE [18] similarity of moment descriptions before and after the substitution is expected to be above a threshold (0.4 in our implementation). The second check is to ensure no person names appear in the substituted version. We detect person names in the substituted moment description using Flair [19]. If a substitution fails to pass both checks, the moment description undergoes human review and is fed to ChatGPT again for another substitution with a different temperature parameter setting, till it passes the two checks.
The above procedure replaces 160,701 words across 72,842 moment descriptions, averaging 2.21 words per description. Table 3 presents the top 12 most frequently replaced words, which shows a diversity of personal pronouns in the descriptions.333 The implementation details and the prompt are detailed in Appendix A
@@ 3.2 Relevance Annotation and Quality Control
Let $m.c$ and $m.v$ represent the caption and the visual content of moment $m$, respectively. To annotate the ground truth moments for a query $q$, we initially retrieve the top-$K$ moment candidates based on their similarity to $q$ by using SimCSE [18], denoted by using $sim(q,m.c)$. In our annotation process, we set $K=20$ for the first batch. This batch of 20 query-moment pairs is then presented to two annotators. [^4] The query $q$ is the same in the batch of 20 query-moment pairs $\langle q,m.v\rangle$. However, during annotation, an annotator is presented with one query-moment pair each time through the annotation interface. Each annotator independently labels the degree of relevance of every query-moment pair $\langle q,m.v\rangle$ purely based on the moment’s visual content, assigning a score from 0 for irrelevant, to 4 for a perfect match.
If the difference between the two relevance scores assigned by two annotators is either 1 or 0, then we considered the two annotators to have reached a consensus. The average relevance score was then rounded up to the nearest whole number as the final score for this query-moment pair. If the two annotators fail to reach a consensus, the same query-moment pair is assigned to another two annotators. Then we have a total of 4 scores. Among the 4 scores, we remove one highest score and one lowest relevance score. If the difference between the remaining two scores is either 1 or 0, we consider a consensus is reached; the average of the remaining two scores is rounded up to the nearest whole number as the final score. Otherwise, the pair is discarded.
@@ 4 Evaluation Metric for RVMR
NDCG for Ranked Retrieval. Normalized Discounted Cumulative Gain (NDCG) is delicately designed for evaluating ranking results with different relevance levels [20]. Specifically, the Discounted Cumulative Gain (DCG) of the top $K$ ranked results is defined in Equation 1, where $i$ is the ranking position with 1 being the top ranked position, $rel_{i}$ is the level of relevance. For example, the left part of Figure 2(b) shows four ground truth moments with $g_{1}$ at rank 1 position and $g_{4}$ at rank 4 position. To their left are the relevance levels with $rel_{1}=4$ for $g_{1}$ and $rel_{2}=2$ for $g_{2}$.
@@
Figure 2 shows the computation of $NDCG@3,IoU\geq 0.3$ as an example. Starting with top predicted moment $p_{1}$, where $IoU(p_{1},g_{1})=0.35$ and $IoU(p_{1},g_{3})=0.4$, both exceeding the threshold $\mu=0.3$, as shown in Figure 2(b). Since the $IoU$ with $g_{3}$ is higher, we consider $p_{1}$ matching with $g_{3}$ and assign $g_{3}$’s relevance score to $p_{1}$. Then $g_{3}$ is removed from the ground truth due to being matched by a predicted moment. Assuming $p_{2}$ is very similar to $p_{1}$ (or a near duplicate), [^5] This is a rare case, and it is unlikely to have two ground truth moments matching duplicate predictions as well. However, we would like to show that our proposed measure is able to handle such a rare case. with $IoU(p_{2},g_{1})=0.35$ and $IoU(p_{2},g_{3})=0.4$, as depicted in Figure 2(c). Since $g_{3}$ has been removed from the ground truth, $g_{1}$ becomes the sole match for $p_{2}$, resulting in a relevance score of 4 for $p_{2}$, and the removal of $g_{1}$ from the ground truth. For $p_{3}$, $IoU(p_{3},g_{4})=0.5$, indicating a match with $g_{4}$, thus $p_{3}$ is assigned a relevance score of 2, and $g_{4}$ is removed. The relevance scores of the predicted moments are 2, 4, and 2. To compute $NDCG$ of the ground truth ranking, we consider a perfect ranking of the top 3 ground truth moments (4, 2, and 2), regardless of whether they are matched by predicted moments or not, as illustrated by $K=3$ on the left side of Figure 2(b).
@@ 5 Baseline Performance
Illustrated in Figure 1, the closest task setting to RVMR is VCMR. In particular, if a VCMR model can compute a form of confidence for its retrieval result, then a ranking of the predicted moments can be easily achieved. Hence, we adapt three VCMR models to RVMR and evaluate them on the TRV-Ranking :XML [6], CONQUER [21], and ReLoCLNet [22]. The main change in the adaptations is the introduction of weight to the training loss based on query-moment similarity i.e., $sim(q,m.c)$, recognizing that moments vary in relevance to a query. This weight aims to diminish the influence of less relevant moments on model training by adjusting the loss. The implementation details are in Appendix C.1.
@@
XML [6] 1 0.1010 0.0923 0.0737 0.0662 0.0258 0.0269
20 0.2331 0.2243 0.1700 0.1650 0.0627 0.0664
40 0.2114 0.2167 0.1530 0.1590 0.0583 0.0635
CONQUER [21] 1 0.0952 0.0835 0.0808 0.0687 0.0526 0.0484
20 0.2130 0.1995 0.1976 0.1867 0.1527 0.1368
40 0.2183 0.1968 0.2022 0.1851 0.1524 0.1365
ReLoCLNet [22] 1 0.1504 0.1439 0.1303 0.1269 0.0866 0.0849
20 0.3815 0.3792 0.3462 0.3427 0.2381 0.2386
40 0.4418 0.4439 0.4060 0.4059 0.2787 0.2877
We conducted three sets of experiments with different $N=\{1,20,40\}$ values for a comprehensive evaluation, where $N$ is the number of moments included in the pseudo training set for each query, by the query-caption similarity $sim(q,m.c)$. We use evaluation metric $NDCG@K,IoU\geq\mu$, for $K=\{10,20,40\}$ and $\mu=\{0.3,0.5,0.7\}$. The parameter search is based on the best results on the validation set with $NDCG@20,IoU\geq 0.5$. Table 5 reports the experiment results with $NDCG@20$, and the full results are in Appendix C.2.
@@
The $K$ Values for $NDCG$. When training with the top 1 pseudo training set, no clear pattern emerges across the three models as $K$ values change. With the top 20 and top 40 pseudo training sets, the $NDCG$ slightly increases as $K$ changes from 10 to 40 for XML and ReLoCLNet, while CONQUER shows the opposite trend.
@@
The Choice of $N$ in Pseudo Training: The $N=1$ training set yields the lowest scores across all three models, suggesting insufficient training instances. For XML, the best performance is achieved with $N=20$, while ReLoCLNet performs best with $N=40$. The results for CONQUER are comparable for $N=20$ and $N=40$. These findings indicate that the models have different capabilities in handling noise in the training data.
Comparison of Baseline Models: All three models (XML, CONQUER, and ReLoCNet) exhibit consistent performance trends across different training sets and metrics. Among the three, ReLoCNet emerges as the top performer, notably when sufficient training moments are provided i.e., $N=40$. However, the three baselines show different performance ranking on the VCMR task, as reported in the original papers [6, 21, 22], where CONQUER demonstrates superior performance against ReLoCNet, and XML is slightly better than ReLoCNet as well. The discrepancy on RVMR implies that the abilities required by our RVMR task differ from those of the VCMR task.